You are standing in line at a ticket window. There are five more people behind you than ahead of you. The total number of people in line (including you) is three times the number of people ahead of you. How many people are ahead of you?
Found 3 solutions by ewatrrr, stanbon, ikleyn:
Answer by ewatrrr(24785)

There are five more people behind you than ahead of you
(x+ 5) + 1 + x = 3x Stated
6 = x
6 people are ahead of you

Answer by stanbon(75887)

You are standing in line at a ticket window. There are five more people behind you than ahead of you. The total number of people in line (including you) is three times the number of people ahead of you. How many people are ahead of you?
----------------
b - a = 5
b + a = 3a
---------------------
Subtract and solve for "a":
2a = 3a-5
a = 5 (# of people ahead of you)
-------------
Cheers,
Stan H.
------------

Answer by ikleyn(52781)

.
You are standing in line at a ticket window. There are five more people behind you than ahead of you. The total number of people in line (including you) is three times the number of people ahead of you. How many people are ahead of you?
~~~~~~~~~~~~~~~~~~~~~~

b - a = 5
b + a + 1 = 3a ("including you")
---------------------
Subtract and solve for "a":
2a+1 = 3a-5
a = 6 (# of people ahead of you)